OpenStudy (anonymous):

inverse of f(x)=2x-4?

OpenStudy (anonymous):

did you get the last one?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

there was an error, so check it. then do this one the same way

OpenStudy (anonymous):

i dont undeerstand!

Parth (parthkohli):

\(\Large \color{midnightblue}{\rightarrow f^{-1}(x) = {x + 4 \over 2} }\) Can anyone explain me too?

OpenStudy (amistre64):

if you did not get the last one, then you need to work with the people who are trying to help you learn it.

Parth (parthkohli):

Am I right?

Parth (parthkohli):

@jabbers! Let's learn together.

OpenStudy (anonymous):

kk

Parth (parthkohli):

1) Add 4 2) Divide by 2.

OpenStudy (anonymous):

k

Parth (parthkohli):

I am not sure, we'll have to check.

OpenStudy (anonymous):

k

OpenStudy (turingtest):

I'll probably just repeat what satellite has said, but... let y=f(x) we then have y=2x-4 solve for x, then switch the places of x and y the new expression for y is the inverse f^(-1)

OpenStudy (anonymous):

easiest method is to write \[y=2x-4\]switch x and y because that is what the inverse function does and write \[x=2y-4\] and then solve for \(y\) via \[x=2y-4\] \[x+4=2y\] \[y=\frac{x+4}{2}\] so \[f^{-1}(x)=\frac{x+4}{2}\]

OpenStudy (anonymous):

ok i understand. thanks!

OpenStudy (anonymous):

eyeball method also works your function says 1) multiply by 2 2) subtract 4 inverse will say 1) add 4 2) divide by 2 you write this as \[f^{-1}(x)=\frac{x+4}{2}\]

Parth (parthkohli):

@satellite73 Why is it another variable?

OpenStudy (anonymous):

not sure what you are asking. we usually write functions as a function of x, but the variable is unimportant, only the action matters. we could write \[f^{-1}(x)=\frac{x+4}{2}\] or \[f^{-1}(y)=\frac{y+4}{2}\] or \[f^{-1}(\zeta)=\frac{\zeta+4}{2}\] or even \[f^{-1}(\spadesuit)=\frac{\spadesuit +4}{2}\]

Parth (parthkohli):

lol

Parth (parthkohli):

Why do we get this equation? ---> 2y + 4 = x

OpenStudy (anonymous):

if i recall correctly indian mathematicians where the first to use variables, and they used colors

Parth (parthkohli):

\(\Large \color{midnightblue}{\rightarrow f^{-1}(x) = f(y) }\) Right?